OpenStudy (anonymous):

George invested some money at 5%. George also invested $180 more than 5 times that amount at 8%. How much is invested at each rate if George receives $1663.65 in interest after one year? (round to two decimal places if necessary.)

OpenStudy (anonymous):

Let the money invested be x Interest = PNR/100 =5x/100 =x/20 Other amt. = 5x+180 I=(5x+180)8/100 =40(x+36)/100 2(x+36)/5 And we know, x/20 + 2(x+36)/5 = 1663.65 x=3665

OpenStudy (anonymous):

How did u get, x+y+180?

OpenStudy (anonymous):

sorry, x=y+180

OpenStudy (anonymous):

sorry, I missed the 5 times